Problem 6-5A (Part Level Submission) You are provided with the following information for Koetteritz Inc. for the month ended(a ) Your answer is partially correct. Try again. Calculate ending inventory, cost of goods sold, gross profit under each ofCalculate gross profit rate under each of the following methods. (1) LIFO. (2) FIFO. (3) Average-cost. (Round answers to 1 deCompare the results for the three cost flow assumptions and answer the following questions. In this period of rising prices,
Answer #1
Cost of Goods Available For Sale
Date Description Quantity Unit cost Amount
June 1 Beginning inventory 39 $40 $1,560
June 4 Purchase 137 45 6,165
June 18 Purchase less return 45 47 2,115
June 28 Purchase 27 51 1,377
Total 248 $11,217

Sales = 108-16*$71+67*$75

Sales = $6,532+5,025 = $11,557

Sales in units = 108-16+67 = 159 units

Ending units = 248 - 159 = 89 units

Average cost per unit = $11,217 / 248 = $45.23 per unit

LIFO:

Ending inventory = $1,560 + 2,250 (50*$45) = $3,810

Cost of goods sold = Cost of goods available for sale - Ending inventory

Cost of goods sold = $11,217 - 3,810 = $7,407

Gross profit = Sales - Cost of goods sold

Gross profit = $11,557 - 7,407 = $4,150

FIFO:

Ending inventory = $1,377+2,115+765(17*$45) = $4,257

Cost of goods sold = $11,217 - 4,257 = $6,960

Gross profit = $11,557 - 6,960 = $4,597

Average-cost

Ending inventory = 89*$45.23 = $4,025

Cost of goods sold = 159*$45.23 = $7,192

Gross profit = $11,557 - 7,192 = $4,365

2.

Gross profit rate:

LIFO = $4,150 / 11,557 *100 = 35.9%

FIFO = $4,597 / 11,557 *100 = 39.8%

Average cost = $4,365 / 11,557 *100 = 37.8%

3.

In this period of rising prices, LIFO gives the highest cost of goods sold and the lowest gross profit. FIFO gives the lowest cost of goods sold and the highest gross profit